OpenStudy (anonymous):

please help me out (: find the x coordinate when you solve x-y=30 -x+4y=15 a.15 b.-5 c.30 d.45

OpenStudy (anonymous):

set the first equation to y so y=x-30 and plug it into your second equation so -x+4(x-30)=15 and simplify -x+4x-120=15 3x=135 x=45

OpenStudy (lalaly):

add both equations .. the x cancels 3y=45 y=15 substitute in any one of the equations x-15=30 x=15+30 = 45

OpenStudy (anonymous):

is this the one you were talking about lalaly?

OpenStudy (lalaly):

no, i meant the one with the even consecutive numbers

OpenStudy (anonymous):

what was the answer?

OpenStudy (lalaly):

they didnt explain it to u right, and the answer was 50

OpenStudy (anonymous):

can you explain how i get 50 please?

OpenStudy (lalaly):

can u post the question again xD without the choices only the question

OpenStudy (anonymous):

why was it deleted...?

OpenStudy (lalaly):

i think it was the sum of three consecutive even numbers equals 156. what is the least of these numbers?